Consider a firm that had been priced using a 12 percent growth rate and a 14 percent required return. The firm recently paid a $1.45 dividend. The firm just announced that because of a new joint venture, it will likely grow at a 12.5 percent rate.

How much should the stock price change (in dollars and percentage)? (Round your answers to 2 decimal places.)
Answer #1

Po - Divi -9 Po = Price of Stock Divi = Estimated Dividends for Next Period r = Required Rate of Return 9 = Growth Rate

a) g = 12%

Div1 = $1.45 * (1 + 12%) = $1.624

1.624 Po = 14-0.12

Po = 81.20

b) g = 12.5%

Div1 = $1.45 * (1 + 12.5%) = $1.63125

1.63125 Po = 0.14-

Po= 108.75

Change in Stock Prices = $108.75 - $81.20 = $27.55

Percentage Change = $27.55/$81.20 = 33.93%